# From Pixels to Titles: Video Game Identification by Screenshots using Convolutional Neural Networks

### v2

@@ 6 Conclusions
This paper explores the application of five distinct CNN architectures (MobileNet, DenseNet, EfficientNetB0, EfficientNetB2, and EfficientNetB3) for identifying video games through screenshots across 22 diverse home console systems, from Atari 2600 (first released in 1977) to PlayStation 5 (first released in 2020). This is a pioneering work as it is the first attempt to identify video game titles by their screenshots. The computer simulations confirmed the hypothesis that CNN’s inherent capacity of automatically extracting relevant features from images is sufficient to identify video game titles from single screenshots in most scenarios, without relying on other features. Using pre-trained weights from the ImageNet dataset, an average accuracy of $74.51\%$ over all the $22$ systems is achieved with the EfficientNetB3 architecture. On the other hand, DenseNet169 architecture is the best in 14 of the 22 tested systems.
@@
When weights pre-trained in another screenshot dataset (Arcade) are used as initial weights - instead of those from ImageNet - the accuracy improves for both EfficientNetB2 and EfficientNetB3, at the same time that the amount of epochs required to converge lowers. The average accuracy over all the $22$ systems increases from $74.51\%$ to $76.36\%$ with EfficientNetB3, while the number of epochs required for models to converge decreases from $23.7$ to $20.5$. On the other hand, the Arcade weights do not improve DenseNet169 in most scenarios.
@@
Overall, considering only the best architecture and weights for each system, an accuracy of $77.67\%$ is achieved. EfficientNetB3 is responsible for the best results in 19 from the 22 systems. The “Arcade weights” are responsible for the best results in 17 of the 22 systems, confirming the hypothesis that starting from weights trained on a different dataset but the same task are better than starting with the more general ImageNet weights in most scenarios.
@@
The simulations performed for this paper showed evidence of the efficacy of CNNs in the task of video game identification by screenshots. Since the largest networks explored in this paper achieved the best results, future research will explore even larger CNN architectures and CNN ensembles to further enhance accuracy. Additionally, this study suggests potential applications in other screenshot-based tasks, such as genre classification or similar game searches, leveraging the efficacy of CNNs in video game identification.

### v3

@@ VIII Conclusions
This paper explores the application of ten distinct CNN architectures (VGG16, ResNet50, ResNet152, MobileNet, DenseNet169, DenseNet201, EfficientNetB0, EfficientNetB2, EfficientNetB3, and EfficientNetV2S) and three transformers architectures (ViT-B16, ViT-L32, and SwinT) for identifying video games through screenshots across 22 diverse home console systems, from Atari 2600 (first released in 1977) to PlayStation 5 (first released in 2020). This is a pioneering work as it is the first attempt to identify video game titles by their screenshots. The experiments confirmed the hypothesis that CNN’s inherent capacity of automatically extracting relevant features from images is sufficient to identify video game titles from single screenshots in most scenarios, without relying on other features. Using pre-trained weights from the ImageNet dataset, an average accuracy of $77.44\%$ over all the $22$ systems is achieved with the EfficientNetV2S architecture. It was also the best architecture for 16 of the 22 systems. While transformer architectures consistently outperform CNNs in image classification tasks, this was not the case for identifying video games from screenshots. Although transformers could learn in most scenarios, their average accuracy was lower than that of all CNNs, except for VGG16.
@@
When weights pre-trained on another screenshot dataset (Arcade) are used as initial weights instead of those from ImageNet, the accuracy improves for both EfficientNetB3 and EfficientNetV2S, while the number of epochs required to converge decreases. The average accuracy across all 22 systems increases from $77.44\%$ to $77.63\%$ with EfficientNetV2S, and the number of epochs required for convergence decreases from $26.9$ to $24.5$. Although the average accuracy is slightly higher, only nine of the 22 systems actually show improvements. In contrast, EfficientNetB3 shows a more significant improvement, with average accuracy increasing from $74.51\%$ to $76.36\%$, and 19 of the 22 systems showing improvements. Additionally, the number of epochs required for convergence decreases from $24.4$ to $19.9$. Interestingly, DenseNet201 shows improvements in 10 of the 22 systems, but the average accuracy remains the same regardless of whether ImageNet or Arcade weights are used. However, the epochs required for convergence still decrease from 24.6 to 22.0.
@@
Overall, considering only the best architecture and weights for each system, an accuracy of $78.79\%$ is achieved. EfficientNetV2S is responsible for the best results in 15 from the 22 systems. The “Arcade weights” are responsible for the best results in 12 of the 22 systems, i.e., starting from weights trained on a different dataset, but the same task, are better than starting with the more general ImageNet weights in most scenarios.
@@
The experiments conducted for this paper provided evidence of the efficacy of CNNs in identifying video games from screenshots. Since the largest EfficientNet architecture explored in this paper achieved the best results, future research will explore even larger CNN architectures and CNN ensembles to further enhance accuracy. The challenges of adding new games to trained networks and identifying a game title without knowing its system a priori should also be investigated, as discussed in Section VII. Additionally, this study suggests potential applications in other screenshot-based tasks, such as genre classification or similar game searches, leveraging the efficacy of CNNs in video game identification.